OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. C = 71°, a = 27, c = 26 A = 79.1°, B = 29.9°, b = 49.3; A = 100.9°, B = 8.1°, b = 49.3 A = 10.9°, B = 98.1°, b = 24.8; A = 169.1°, B = 81.9°, b = 24.8 A = 79.1°, B = 29.9°, b = 13.7; A = 100.9°, B = 8.1°, b = 3.9 A = 10.9°, B = 98.1°, b = 27.2; A = 169.1°, B = 81.9°, b = 27.2

OpenStudy (anonymous):

Law of sines: \[\frac{\sin A}{a}=\frac{\sin B}{b}=\frac{sin C}{c}=K\] K = some const. 1 use C and c to find K, 2 use K, and a to find the two values of A

OpenStudy (anonymous):

Not gonna lie, I don't feel like doing the work...

OpenStudy (snuggielad):

now that i have an aswer to....you have to do the work...if you dont you dont get help...refer to the code of conduct...nobody here is dumb enough to give you the answers

mathslover (mathslover):

I think you have to use the law of sines as in the question it is given that "use the law of sines to solve the triangles"

OpenStudy (snuggielad):

but your honest...thats good...you still need to get help...but you have to come to your own answers based on what you learn